OpenStudy (anonymous):

3d^9 X 2d^-2? simplify

OpenStudy (anonymous):

if your multiplying polynomials you need to multiply and then add the exponents like this 3 X 2 = 6 d^9 + d^ -2 = d^7 6d^7

OpenStudy (anonymous):

whats the answer

OpenStudy (goldphenoix):

\[\large \large 3d^9 \times 2d^-2\] Since they have different base, you would want to multiply the base. So what does 3 *2 give you?

OpenStudy (goldphenoix):

Also, you would also want to add the exponent. So 9 + (-2) would give you?

OpenStudy (anonymous):

11

OpenStudy (goldphenoix):

When your adding a number by a negative number, then it would become a negative.